Graph an ellipse with foci at (0, ±2 times square root of 10 ) and x-intercept at (9, 0).
Answer by lwsshak3(11628)

Graph an ellipse with foci at (0, ±2 times square root of 10 ) and x-intercept at (9, 0)
**
Standard form of equation of an ellipse with vertical major axis:
(x-h)^2/b^2+(y-k)^2/a^2=1, a>b, (h,k)=(x,y) coordinates of center
**
For given ellipse:
center:(0,0)
c=1/2 distance between foci=2√10
c^2=40
..
b=9
b^2=81
..
c^2=a^2-b^2
a^2=c^2+b^2=40+81=121
a=√121=11
..
Equation:
x^2/81+y^2/121=1
see graph below as a visual check:
y=±(121-121x^2/81)^.5